OpenStudy (anonymous):

use mathematical induction to prove that the statement is true for every positive integer n 1+3+5+...+(2n-1)= n^2

OpenStudy (anonymous):

add the next term and show that the sum equals (n+1)^2

OpenStudy (dumbcow):

next term would be (2n+1) --> n^2 + 2n +1 = (n+1)^2

OpenStudy (dumbcow):

also show initial step n=1 is true 1 = 1^2

OpenStudy (anonymous):

thank you guys